# Distinguishing Neighborhood Representations Through Reverse Process of GNNs for Heterophilic Graphs

### v1

@@ A.3 Forward Pass of Reverse Process
To calculate Equation 9, the initial step involves computing $(-1)^{k}\hat{\mathbf{A}}^{k}\mathbf{X}^{(\ell)}\mathbf{W}^{k}$. The time complexity for this part is $O(k|\mathcal{E}|d+|\mathcal{V}|d^{2}+kd^{3})$. Summing over $k$ from 0 to $K$, the overall complexity becomes $O(K^{2}|\mathcal{E}|d+K|\mathcal{V}|d^{2}+K^{2}d^{3})$. With $L_{R}$ representing the number of reverse process layers, the total time complexity is calculated as $O(K^{2}L_{R}|\mathcal{E}|d+KL_{R}|\mathcal{V}|d^{2}+K^{2}L_{R}d^{3})$.
@@ A.4 Backward Pass of Reverse Process
The time complexity of the backward pass primarily hinges on the computation of $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$. By applying the chain rule, this can be expressed as:
@@
Given $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}$, $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell+1)}}$ can be derived as follows:
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell+1)}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\frac{\partial\mathbf{X}^{(\ell)}}{\partial\mathbf{X}^{(\ell+1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\frac{\partial}{\partial\mathbf{X}^{(\ell+1)}}\left(\mathbf{X}^{(\ell+1)}+\sum_{k=1}^{K}(-1)^{k}\hat{\mathbf{A}}^{k}\mathbf{X}^{(\ell+1)}\mathbf{W}^{k}\right)$
$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}+\sum_{k=1}^{K}(-1)^{k}\left(\hat{\mathbf{A}}^{k}\right)^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(\ell)}}\left(\mathbf{W}^{k}\right)^{\top}.$ (16)
And we can compute $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R}+1)}}$ as follows:
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R}+1)}}=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial\mathbf{X}^{(-L_{R})}}{\partial\mathbf{X}^{(-L_{R}+1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial}{\partial\mathbf{X}^{(-L_{R}+1)}}\left(\mathbf{X}^{(-L_{R}+1)}+\sum_{k=1}^{K}(-1)^{k}\hat{\mathbf{A}}^{k}\mathbf{X}^{(-L_{R}+1)}\mathbf{W}^{k}\right)$
$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}+\sum_{k=1}^{K}(-1)^{k}\left(\hat{\mathbf{A}}^{k}\right)^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\left(\mathbf{W}^{k}\right)^{\top}.$ (17)
By recursively applying Equation 16 to Equation 17, we can obtain $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-1)}}$ as follows:
@@
$=$ $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}+\sum_{\ell=2}^{L_{R}-1}\binom{L_{R}-1}{\ell}\sum_{(k_{1},k_{2},\cdots,k_{\ell-1})}(-1)^{\sum_{i=1}^{\ell-1}k_{i}}({\hat{\mathbf{A}}^{\top}})^{\sum_{i=1}^{\ell-1}k_{i}}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)({\mathbf{W}^{\top}})^{\sum_{i=1}^{\ell-1}k_{i}}.$ (18)
In addition, $\frac{\partial\mathbf{X}^{(-1)}}{\partial\mathbf{W}}$ can be obtained as follows:
@@
$\frac{\partial\mathbf{X}^{(-1)}}{\partial\mathbf{W}}$ $=\frac{\partial}{\partial\mathbf{W}}\left(\mathbf{X}^{(0)}+\sum_{k_{0}=1}^{K}(-1)^{k_{0}}\hat{\mathbf{A}}^{k_{0}}\mathbf{X}^{(0)}\mathbf{W}^{k_{0}}\right)$
$=\sum_{k_{0}=1}^{K}(-1)^{k_{0}}k_{0}\frac{\partial}{\partial\mathbf{W}^{k_{0}}}(\hat{\mathbf{A}}^{k_{0}}\mathbf{X}^{(0)}\mathbf{W}^{k_{0}})\mathbf{W}^{k_{0}-1}\frac{\partial\mathbf{W}}{\partial\mathbf{W}}.$ (19)
Finally, $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ can be derived as the multiplication of $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-1)}}$ and $\frac{\partial\mathbf{X}^{(-1)}}{\partial\mathbf{W}}$:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ $=\sum_{k_{0}=1}^{K}\Bigg{[}(-1)^{k_{0}}k_{0}(\hat{\mathbf{A}}^{k_{0}}\mathbf{X}^{(0)})^{\top}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)\mathbf{W}^{k_{0}-1}$
@@
$+\sum_{\ell=2}^{L_{R}-1}\binom{L_{R}-1}{\ell}\sum_{(k_{1},k_{2},\cdots,k_{\ell-1})}(-1)^{\sum_{i=1}^{\ell-1}k_{i}}k_{0}(\hat{\mathbf{A}}^{k_{0}}\mathbf{X}^{(0)})^{\top}({\hat{\mathbf{A}}^{\top}})^{\sum_{i=1}^{\ell-1}k_{i}}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)({\mathbf{W}^{\top}})^{\sum_{i=1}^{\ell-1}k_{i}}\mathbf{W}^{k_{0}-1}\Bigg{]}.$ (20)
@@
By the same procedure in the forward process, the time complexity of the first term in brackets is $O(k_{0}(|\mathcal{E}|d+d^{3})+|\mathcal{V}|d^{2})$. In the case of the second term, there are up to $2\ell K$ possible outcomes for $\sum_{i=1}^{\ell}k_{i}$. So the time complexity is $O(L_{R}+K\ell|\mathcal{V}|d^{2}+\ell^{2}K^{2}(d^{3}+|\mathcal{E}|d))$. So the overall time complexity of reverse process is $O((L_{D}^{2}+M^{2}L_{R}^{3})|\mathcal{E}|d+(L_{D}^{2}+ML_{R}^{2})|\mathcal{V}|d^{2}+M^{2}L_{R}^{3}d^{3})$. The memory complexity is $O(|\mathcal{V}|d+|\mathcal{E}|+d^{2})$.
@@ Appendix C Multi-head Attention for GAT
In this case, the upper bound of the Lipschitz constant of $h$ is calculated as follows:
$\text{Lip}(h)$ $=\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\frac{1}{K}\sum_{k=1}^{K}\hat{\mathbf{A}}^{(k)}\mathbf{X}^{(\ell)}\mathbf{W}^{(k)}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}$ (23)
$\leq\frac{1}{K}\sum_{k=1}^{K}\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\hat{\mathbf{A}}^{(k)}\mathbf{X}^{(\ell)}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}\lVert{\mathbf{W}^{(k)}}\rVert_{F}$ (24)

### v2

@@ A.3 Forward Pass of Reverse Process
To calculate Equation 11, the initial step involves computing $(-1)^{m}\hat{\mathbf{A}}^{m}\mathbf{X}^{(\ell)}\mathbf{W}^{m}$. The time complexity for this part is $O(m|\mathcal{E}|d+|\mathcal{V}|d^{2}+md^{3})$. Summing over $m$ from 0 to $M$, the overall complexity becomes $O(M^{2}|\mathcal{E}|d+M|\mathcal{V}|d^{2}+M^{2}d^{3})$. With $L_{R}$ representing the number of reverse process layers, additional time complexity becomes $O(M^{2}L_{R}|\mathcal{E}|d+ML_{R}|\mathcal{V}|d^{2}+M^{2}L_{R}d^{3})$ comparing to forward pass without reverse process.
@@ A.4 Backward Pass of Reverse Process
The time complexity of the backward pass is primarily determined by the computation cost of $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$. By applying the chain rule, the gradient can be expressed as:
@@
By sequentially multiplying $\frac{\partial\mathbf{X}^{(-n)}}{\partial\mathbf{X}^{(-n+1)}}$ for $n=L_{R},\cdots,2$ to the right side of $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}$, we derive:
@@
$=$ $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}+\sum_{\ell=1}^{L_{R}-1}\binom{L_{R}-1}{\ell}\sum_{(m_{1},m_{2},\cdots,m_{\ell})}(-1)^{\sum_{i=1}^{\ell}m_{i}}({\hat{\mathbf{A}}^{\top}})^{\sum_{i=1}^{\ell}m_{i}}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)({\mathbf{W}^{\top}})^{\sum_{i=1}^{\ell}m_{i}}\;,$ (18)
where $m_{i}=1,\cdots,M$ for all $i$. To demonstrate Equation 18, we show part of the sequential multiplication process. Multiplying $\frac{\partial\mathbf{X}^{(-L_{R})}}{\partial\mathbf{X}^{(-L_{R}+1)}}$ to the right side of $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}$, we get:
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial\mathbf{X}^{(-L_{R})}}{\partial\mathbf{X}^{(-L_{R}+1)}}$ $=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial}{\partial\mathbf{X}^{(-L_{R}+1)}}\left(\mathbf{X}^{(-L_{R}+1)}+\sum_{m=1}^{M}(-1)^{m}\hat{\mathbf{A}}^{m}\mathbf{X}^{(-L_{R}+1)}\mathbf{W}^{m}\right)$
$=\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}+\sum_{m=1}^{M}(-1)^{m}\left(\hat{\mathbf{A}}^{m}\right)^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\left(\mathbf{W}^{m}\right)^{\top}\;,$ (19)
which can also be obtained by substitute $L_{R}=2$ in Equation 18.
@@
Next, we multiply $\frac{\partial\mathbf{X}^{(-L_{R}+1)}}{\partial\mathbf{X}^{(-L_{R}+2)}}$ on the right side of Equation 19, yielding:
$\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\frac{\partial\mathbf{X}^{(-L_{R})}}{\partial\mathbf{X}^{(-L_{R}+1)}}\frac{\partial\mathbf{X}^{(-L_{R}+1)}}{\partial\mathbf{X}^{(-L_{R}+2)}}$
$=$ $\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}+\sum_{m=1}^{M}(-1)^{m}\left(\hat{\mathbf{A}}^{m}\right)^{\top}\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\left(\mathbf{W}^{m}\right)^{\top}\right)\frac{\partial}{\partial\mathbf{X}^{(-L_{R}+2)}}\left(\mathbf{X}^{(-L_{R}+2)}+\sum_{m=1}^{M}(-1)^{m}\hat{\mathbf{A}}^{m}\mathbf{X}^{(-L_{R}+2)}\mathbf{W}^{m}\right)$
@@
$=$ $\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}+\sum_{\ell=1}^{2}\binom{2}{\ell}\sum_{(m_{1},m_{2},\cdots,m_{\ell})}(-1)^{\sum_{i=1}^{\ell}m_{i}}({\hat{\mathbf{A}}^{\top}})^{\sum_{i=1}^{\ell}m_{i}}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)({\mathbf{W}^{\top}})^{\sum_{i=1}^{\ell}m_{i}}.$ (20)
Repeating the process above, we can derive Equation 18.
Finally, $\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ is derived by multiplying Equation 18 with $\frac{\partial\mathbf{X}^{(-1)}}{\partial\mathbf{W}}$:
$\frac{\partial\mathcal{L}}{\partial\mathbf{W}}$ $=\sum_{m_{0}=1}^{M}(-1)^{m_{0}}m_{0}(\hat{\mathbf{A}}^{m_{0}}\mathbf{X}^{(0)})^{\top}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)\mathbf{W}^{m_{0}-1}$
$+\sum_{\ell=1}^{L_{R}-1}\binom{L_{R}-1}{\ell}\sum_{(m_{0},m_{1},\cdots,m_{\ell})}(-1)^{\sum_{i=1}^{\ell}m_{i}}m_{0}(\hat{\mathbf{A}}^{m_{0}}\mathbf{X}^{(0)})^{\top}({\hat{\mathbf{A}}^{\top}})^{\sum_{i=1}^{\ell}m_{i}}\left(\frac{\partial\mathcal{L}}{\partial\mathbf{X}^{(-L_{R})}}\right)({\mathbf{W}^{\top}})^{\sum_{i=1}^{\ell}m_{i}}\mathbf{W}^{m_{0}-1}\;,$ (21)
@@
where $m_{i}=1,\cdots,M$ for all $i$. The time complexity of the first term is $O(M^{2}|\mathcal{E}|d+M|\mathcal{V}|d^{2}+M^{2}d^{3})$. In the case of the second term, the time complexity is $O(M^{2}L_{R}^{3}|\mathcal{E}|d+ML_{R}^{2}|\mathcal{V}|d^{2}+M^{2}L_{R}^{3}d^{3})$, since there are up to $2\ell M$ possible outcomes for $\sum_{i=1}^{\ell}m_{i}$. Therefore, the overall time complexity of the reverse process is $O((L_{F}^{2}+M^{2}L_{R}^{3})|\mathcal{E}|d+(L_{F}^{2}+ML_{R}^{2})|\mathcal{V}|d^{2}+M^{2}L_{R}^{3}d^{3})$. The memory complexity is $O(|\mathcal{V}|d+|\mathcal{E}|+d^{2})$.
@@ Appendix C Multi-Head Attention for GAT
In this case, the Lipschitz constant of $h$, $\text{Lip}(h)<1$ is satisfied if
$\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\frac{1}{K}\sum_{k=1}^{K}\hat{\mathbf{A}}^{(k)}\mathbf{X}^{(\ell)}\mathbf{W}^{(k)}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}<1\;.$ (24)
The upper bound of left side is computed by:
$\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\frac{1}{K}\sum_{k=1}^{K}\hat{\mathbf{A}}^{(k)}\mathbf{X}^{(\ell)}\mathbf{W}^{(k)}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}$ $\leq\frac{1}{K}\sum_{k=1}^{K}\sup_{\mathbf{X}\neq\mathbf{0}}\frac{\lVert{\hat{\mathbf{A}}^{(k)}\mathbf{X}^{(\ell)}}\rVert_{F}}{\lVert{\mathbf{X}}\rVert_{F}}\lVert{\mathbf{W}^{(k)}}\rVert_{F}$